# Scaling Laws for Task-Optimized Models of the Primate Visual Ventral Stream

### v1

@@ Appendix A Validation on private data
Appendix B Pretrained Models
As part of our comprehensive evaluation, we benchmarked a diverse set of pretrained models sourced from both PyTorch’s native library and the timm library. We tested a total of 121, including ViT Dosovitskiy et al. (2021), DaViT Ding et al. (2022), LeViT Graham et al. (2021), ConvNeXt Liu et al. (2022a), MobileViT Mehta & Rastegari (2022), MaxVit Tu et al. (2022), FastViT Vasu et al. (2023), and EfficientViT Cai et al. (2022). Each model has varying parameter counts, training sample size, training dataset source, and training objective.
(a) Average alignment
(b) Neural benchmarks
@@ Appendix B Pretrained Models
Appendix C Training Evolution
We observed that the alignment scaling behavior varies significantly across different model architectures, prompting an investigation into their training dynamics to understand the underlying causes. As illustrated in Fig. 9, while the average alignment across various models eventually converges to similar levels with sufficient training, fully convolutional architectures—such as ResNets and EfficientNets—exhibit substantially higher alignment scores at the very beginning of training.
Fig. 10 reveals that this early advantage is primarily due to these models having high neural alignment at initialization, even before any learning from data occurs. This suggests that the architectural features inherent to fully convolutional networks endow them with representations that are already closely aligned with neural data from the primate ventral visual stream.
Further emphasizing this point, Fig. 11 demonstrates that strong inductive biases present in these architectures are key determinants of a model’s initial alignment. These findings underscore that while different architectures may ultimately reach similar levels of alignment after extensive training and scaling, their initial conditions and the efficiency with which they achieve alignment can vary greatly.
Figure 9: Evolution of brain alignment throughout training
Figure 10: Evolution of per region alignment throughout training
Figure 11: Trained vs random models
Appendix D Effect of Training Objective
We are exploring alternative training strategies that may lead to better brain alignment than supervised training alone. Our ongoing experiments indicate that models trained on the same datasets using unsupervised objectives for extended durations can achieve higher alignment scores. Another promising direction is adversarial fine-tuning; we have found that even a weak form of adversarial fine-tuning using the Fast Gradient Sign Method can enhance neural alignment, especially in lower visual regions. These approaches suggest that incorporating unsupervised learning objectives and adversarial training techniques may be effective in further improving the alignment between artificial neural networks and the primate visual system.
(a) Comparison of training objectives
(b) ResNet50 Alignment - Regions
@@ Appendix D Effect of Training Objective
(a) Brain alignment evolution across checkpoints
(b) Per region alignment evolution
Figure 13: Alignment trajectories during training for different training objectives
(a) SimCLR training of ResNet models
(b) Adversarial finetuning of ResNet models
Figure 14: ResNet alignment scaling across different training paradigms
Appendix E Region-wise scatter plots
Figure 15: Region vs Region
Figure 16: Region vs Region - KDE fits

### v2

@@ Appendix C Validation on private data
Appendix D Pretrained Models
As part of our comprehensive evaluation, we benchmarked a diverse set of pretrained models sourced from both torchvision (maintainers & contributors, 2016) and the timm (Wightman, 2019) libraries. We tested a total of 94, including ViT Dosovitskiy et al. (2021), DaViT Ding et al. (2022), LeViT Graham et al. (2021), ConvNeXt Liu et al. (2022a), MobileViT Mehta & Rastegari (2022), MaxVit Tu et al. (2022), FastViT Vasu et al. (2023). Each model varies in parameter count, training sample size, dataset source, and training objective, providing a broad spectrum for analysis.
To verify the generalizability of our findings, we conducted evaluations with these pretrained models, including larger networks like CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023), which are pretrained on richer and more diverse datasets such as LAION Schuhmann et al. (2021; 2022). We also compared variations of these models by examining base pretrained models alongside their fine-tuned counterparts on ImageNet, aiming to investigate the impact of fine-tuning on scaling behavior.
Our results indicate that models with extensive pretraining achieve enhanced behavioral alignment, likely due to their exposure to richer and more varied data. However, similar to models trained solely on ImageNet or EcoSet, these pretrained models still exhibit a saturation effect in neural alignment with the primate visual ventral stream (VVS). This suggests that while larger and more diverse datasets improve behavioral predictability, they do not substantially extend the scaling of neural alignment beyond the observed plateau.
The curves in Figure 9 closely follow the scaling patterns estimated for our trained models shown in Figure 2.c, further validating that the observed saturation is consistent across different pretraining regimes and dataset scales. This reinforces our conclusion that scaling alone is insufficient to overcome the limitations in neural alignment and highlights the need for alternative approaches to improve alignment with neural representations.
@@ Appendix D Pretrained Models
Appendix E Training Evolution
Figure 10: Evolution of per-region alignment throughout training. Models with stronger priors—such as ResNet and EfficientNet—exhibit higher neural alignment initially. However, the gap in representational power diminishes as more generalist models are trained on data.
Figure 11: Some untrained models demonstrate non-zero neural alignment at initialization. Nevertheless, all models start with almost zero behavioral alignment, indicating that initial neural alignment arises from architectural biases rather than learned behavior.
Appendix F Effect of Training Objective
@@
Figure 13: Scaling of SimCLR training across regions
Appendix G Region-wise scatter plots
Figure 14: Region vs. Region Comparisons: This figure shows how the alignment scores for each brain region correlate with those of other regions. The diagonal plots illustrate the relationship between the alignment score of each region and the validation accuracy on ImageNet and EcoSet datasets.